OpenStudy (anonymous):

ordered pair for -3x+y=3

OpenStudy (imstuck):

An ordered pair requires you to solve for x and y and put it into the form (x, y). You could choose any value of x and then solve it for y. Let's try x = 1. -3(1) = -3. -3 + y = 3. Add 3 to both sides t get y = 6. That's one solution, (1, 6). You could do it for any value of x. If x = 2, then 2(-3) = -6. -6 + y = 3. Add 6 to both sides to get y = 9. That's another ordered pair, (2, 9). Get it?

OpenStudy (texaschic101):

you realize there is more then 1 answer

OpenStudy (texaschic101):

unless this is multiple choice